OpenStudy (anonymous):

Determine the slope between the points (-1, -5) and (5, 4)

OpenStudy (anonymous):

(4-(-5))/(5-(-1))

OpenStudy (anonymous):

4-(-5)=m(5-(-1) set up

OpenStudy (anonymous):

solve for m

OpenStudy (anonymous):

I have no idea on how to solve this

OpenStudy (anonymous):

use point slope form y=4 and y1=-5 and x=5, x1=-1 so y1-y=m(x1-x)

OpenStudy (anonymous):

actually its y-y1=m(x-x1)

OpenStudy (anonymous):

m=9/6 or 3/2

OpenStudy (anonymous):

Thank you! I actually solved it myself! (: